Suppose small aircraft arrive at a certain airport according to a Poisson process with rate α = 8 per hour, so that the number of arrivals during a time period of t hours is a Poisson rv with parameter μ = 8t. (Round your answers to three decimal places.) (a) What is the probability that exactly 6 small aircraft arrive during a 1-hour period? What is the probability that at least 14 small aircraft arrive during a 1-hour period? What is the expected value and standard deviation of the number of small aircraft that arrive during a 75-min period? expected value standard deviation (b) (c) What is the probability that at least 25 small aircraft arrive during a 2.5-hour period? What is the probability that at most 18 small aircraft arrive during a 2.5-hour period?
Answer #1

a)

probability(exactly 6) = P(X=6)= {ex/x!}= 0.122
probability(at least 6) = P(X>=6)= 1-P(X<=5)= 1-∑x=0x-1   {ex/x!}= 0.809
probability(at least 14) = P(X>=14)= 1-P(X<=13)= 1-∑x=0x-1   {ex/x!}= 0.034

b)expcted value =8*75/60=10

standard deviation =sqrt(10) =3.162

c)

probability (at least 25)= P(X>=25)= 1-P(X<=24)= 1-∑x=0x-1   {ex/x!}= 0.157
probability (at most 18)= P(X<=18)= x=0x   {ex/x!}= 0.381